OpenStudy (anonymous):

(sqrt 6-sqrt 7)^2multiply

OpenStudy (whpalmer4):

\[(\sqrt{6}-\sqrt{7})^2 = (\sqrt{6}-\sqrt{7})(\sqrt{6}-\sqrt{7})\] Do this exactly like you would \[(a-b)(a-b)\]When you've completed the multiplication, then you can simplify some of the results.